OpenStudy (anonymous):

Which of the following numbers cannot be expressed as the difference of two squares of the intergers? 314159265 314159266 314159267 314159268

OpenStudy (apoorvk):

The third option. Differences of squares of integers cannot end with the digit "7". Squares can end only with digits 0,1,4,5,6,9. (check it out) Now, no way by deducting the units digit of two squares would we get the unit's digit '7'. there you go!

OpenStudy (anonymous):

i think you misunderstand the question

OpenStudy (apoorvk):

Not again??? what does it mean?

OpenStudy (apoorvk):

you mean that if "a" and "b" are two integers, the |a^2-b^2| {...magnitude} can which of the given options. right?

OpenStudy (anonymous):

example 63 = 8^2 - 1^2

OpenStudy (apoorvk):

yeah thats what i did. what's the answer then?

OpenStudy (apoorvk):

ofcourse it isnt possible to find the actual nos. we can only find out which is and which isn't.

OpenStudy (apoorvk):

i mean not possible *practically

OpenStudy (anonymous):

the 3rd option can be expressed as the difference of two squares of the intergers

OpenStudy (apoorvk):

What's the reason then? I 'll bet my bottom rupee 314159267 can not be.

OpenStudy (anonymous):

it can be apoorvk

OpenStudy (apoorvk):

how? state your logic.

OpenStudy (anonymous):

157079634^2 - 157079633^2 = 314159267

OpenStudy (apoorvk):

My big bad again. *poker face* *palm face*. You are a genius alright! so how do you solve this then? Whats the funda behind this?

OpenStudy (apoorvk):

and yeah my logic had a loophole. :/

OpenStudy (anonymous):

no im not genius lol

OpenStudy (apoorvk):

And yeah, I learnt never ever to bet my bottom buck ever. :| so what's the logic then?

OpenStudy (anonymous):

Notice that \[2 n+1=(n+1)^2-n^2 \] Then every odd number is the difference of two squares

OpenStudy (anonymous):

n = ((n + 1) / 2 ) ^2 - ((n - 1) / 2 ) ^2

OpenStudy (anonymous):

314159267 = ((314159267 + 1)/2)^2 - ((314159267 - 1)/2)^2

OpenStudy (apoorvk):

amazing. but then, how do you sort out option B and D?

OpenStudy (apoorvk):

I am not sure if I understand modulo, but I 'll read up on it. Thanks once again! :)

OpenStudy (anonymous):

All integers congruent to 2 (modulo 4) cannot be expressed as the dierence of perfect squares. Only numbers that can be expressed as the dierence of perfect squares are those congruent to 0, 1 or 3 (modulo 4). 314159265= 3141592 x 100 + 65 ≡ 65 ≡1 (mod 4) 314159266= 3141592 x 100 + 66 ≡ 66 ≡2 (mod 4) 314159267= 3141592 x 100 + 67 ≡ 67 ≡3 (mod 4) 314159268= 3141592 x 100 + 68 ≡68 ≡0 (mod 4) 314159269= 3141592 x 100 + 69 ≡69 ≡1 (mod 4) So, of the ve numbers given, only 314159266 cannot be expressed as the dierence of perfect squares whereas the others can be.